OpenStudy (hitaro9):

Prove that 81|(10^(n+1) -9n -10) for every non negative integer n value.

OpenStudy (hitaro9):

So I'm guessing this is a proof by induction problem. I showed the base case of zero, but for (n+1) we have 81| 100(10^n) -9n -19 And I'm not sure how to set it up so that it's clear 81 divides that.

OpenStudy (perl):

ok you want to use induction, thats probably easiest

OpenStudy (hitaro9):

If there's some other way I'm open to it, but the way the problem was written it seems like that's what'd be easiest.

OpenStudy (hitaro9):

Oh no, the original case was 81|(10^(k+1) -9k-10)

OpenStudy (perl):

ok i see

OpenStudy (hitaro9):

The goal is going to be 81| (10^(k+2) -9(k+1)_10)

OpenStudy (perl):

correction* assume that : 81 | 10^(k+1) -9k -19 <many steps later> it is true that: 81 | 10^(k+2) -9(k+1) -19

OpenStudy (hitaro9):

-10 but yeah haha

OpenStudy (hitaro9):

So any ideas how to manipulate it so it's clear 81 divides it?

OpenStudy (hitaro9):

Haha it's fine. I've done the same thing plenty of times.

OpenStudy (perl):

:)

OpenStudy (perl):

well now that you understand the problem, you are 3 quarters there

OpenStudy (owlfred):

@perl it should be 10^(k+2) - 9(k+1) - 10.

ganeshie8 (ganeshie8):

\[10^{n+1} -9n-10 = (1+9)^{n+1}-9n-10 = 1+9(n+1)+ 9^2M - 9n-10 = 9^2M\]

OpenStudy (owlfred):

OK, I found another way to prove this...

OpenStudy (perl):

$$ 10^{n+1} -9n-10 = (1+9)^{n+1}-9n-10 \\= 1+9(n+1)+ 9^2M - 9n-10 = 9^2M$$

ganeshie8 (ganeshie8):

yes :)

OpenStudy (owlfred):

Assume that \(10^{k+1} - 9k - 10\) is divisible by 81.\[n = k+1:\]\[10\cdot 10^{k+1} - 9(k + 1) - 10\]\[= 10\cdot 10^{k+1} - 9k - 19\]\[= 10\cdot 10^{k+1} - 90k - 100 + 81 k - 81 \]\[= 10(10^{k+1} - 9 k - 10) + 81 (k - 1)\]

OpenStudy (hitaro9):

Okay. I think I'm following what you did Ganeshie. I get what to do now, thank you.

ganeshie8 (ganeshie8):

These come in handy when you seek an alternative for induction proofs like these : ``` (1+x)^n = 1+xn + x^2(stuff) (1+x)^n = 1+x(stuff) ```

OpenStudy (perl):

if you use ganeshie's argument in your inductive step, but then you won't need to cite the inductive hypothesis

OpenStudy (hitaro9):

Alright then. I never would've seen something like that on my own, thanks a ton.

OpenStudy (perl):

ganeshie's argument is a constructive approach, you are actually showing how to produce a factor of 81 for any nonnegative integer of the form 10^(n+1) -9n -10

OpenStudy (perl):

you can solve for k in terms of n . 10^(n+1) -9n -10 = 18k

OpenStudy (perl):

so it stands as a proof by itself

OpenStudy (hitaro9):

Right. It's pretty creative. I'm not sure how I would think of something like that on an exam

OpenStudy (hitaro9):

But it was super helpful here, thanks a ton.

ganeshie8 (ganeshie8):

I don't think it is easy to think of it on one's own. One needs others to show that it can be done like that... Once you see it can be done like that, it becomes easy to apply it in exam

ganeshie8 (ganeshie8):

Induction proofs are awesome because you always know what do do : where to start and how to proceed etc..

OpenStudy (owlfred):

Induction works almost every time. Just try to express the \(n+1\)-case in terms of the \(n\)-case, and you'll see how things begin to converge.

OpenStudy (hitaro9):

Right. A lot of the other problems have been pretty straight forward. This was the first one that required creativity like that. Again, thanks for pointing out the method.